(2x+3)(5x-7)=(10x+1)(x+6)

Answer by checkley71(8403)

(2x+3)(5x-7)=(10x+1)(x+6)
10x^2+x-21=10x^2+61x+6
x-61x=6+21
-60x=27
x=27/-60
x=-9/20 or .45 answer.
proof:
(2*-.45+3)(5*-.45-7)=(10*-.45+1)(-.45+6)
(-.9+3)(-2.25-7)=(-4.5+1)(5.55)
-2.1*-9.25=3.5*5.55
19.425=19.425